# On the Invariant and Geometric Structure of the Holomorphic Unified Field Theory

### v1

@@ 6 Quantum Mechanics and Quantum Field Theory from HUFT
We define the kinetic momentum operator $\hat{\Pi}_{\mu}:=-i\hbar D_{\mu}$, the position operator $\hat{X}^{\mu}$, multiplication by the coordinate function on $\Sigma_{t}$, and for any bundle endomorphism observable $\hat{O}$ define the expectation value on $\Sigma_{t}$ by:
@@
We take dynamics to be generated by the minimally coupled covariant Hamiltonian $\hat{H}$ either Dirac, Pauli, or Schrödinger built from $h$ and $D_{\mu}$, so that $i\hbar\,\partial_{t}\psi=\hat{H}\psi$ and the probability current is conserved.
We assume standard falloff, so boundary terms at spatial infinity vanish, and take $V$ a real scalar section such as a gauge singlet potential. Then the expectation values of position and kinetic momentum obey:
@@
Conservation of probability current gives the covariant continuity equation $\partial_{t}(\sqrt{\gamma}\,\psi^{\dagger}\psi)+\nabla_{i}(\sqrt{\gamma}\,j^{i})=0$, so boundary terms from $\partial_{t}\sqrt{\gamma}$ and $\nabla_{i}$ cancel in the time derivative of (110). Thus, we obtain the Heisenberg identity in covariant form:
@@
with the Levi–Civita correction precisely reproducing the $\Gamma$–terms below.
For $\hat{O}=\hat{X}^{\mu}$, minimal coupling implies $\hat{\Pi}_{\alpha},\hat{X}^{\mu}]=-i\hbar\,\delta^{\mu}{}_{\alpha}$, and for the kinetic Hamiltonian $\hat{H}_{\rm kin}=\tfrac{1}{2m}\,h^{\alpha\beta}\hat{\Pi}_{\alpha}\hat{\Pi}_{\beta}$ we obtain:
@@
yielding (111) after taking expectation values and accounting for the foliation, no explicit time dependence of $\hat{X}^{\mu}$.
For $\hat{O}=\hat{\Pi}_{\mu}$ we use curvature and metric compatibility and define the kinetic momentum operator:
@@
meaning $h^{\alpha\beta}$ is covariantly constant, so it commutes with $\hat{\Pi}_{\mu}$ so, no extra terms from derivatives of the metric when you commute things. Then the Heisenberg force operator follows:
@@
the right-hand side is the symmetrized product of momentum with field strength, so the operator is Hermitian. Whose symmetric part gives the Lorentz-force term $q\,F_{\mu\nu}\,\dot{X}^{\nu}$ in expectation values. The scalar potential contributes $(i/\hbar)[V,\hat{\Pi}_{\mu}]=-\nabla_{\mu}V$. Finally, when transporting the expectation value with the time-dependent volume form and frame, the Levi–Civita piece from $D/dt$ supplies $\langle\Gamma^{\rho}_{\mu\nu}\,\hat{\Pi}_{\rho}\,\dot{X}^{\nu}\rangle$, completing (112). Since on the real slice $B_{\mu\nu}=F_{\mu\nu}$, the gauge-force term is literally the antisymmetric part of the Hermitian metric.
In a local inertial frame at a point $\Gamma=0$ the relations reduce to by taking expectation values and writing:
@@
When transporting the expectation value, that’s the gravitational part. That extra term is what we call the Levi–Civita piece. It’s the correction needed so that the equation of motion is tensorial coordinate-independent. Without it, you’d be differentiating components as if the basis were fixed, flat space. With it, the quantum expectation values reduce to the geodesic with Lorentz force in the classical limit.
@@
On a static slice of flat space with coordinates $(t,\mathbf{x})$, $\Gamma=0$ and $h_{\mu\nu}\to\eta_{\mu\nu}$. Writing $\hat{\bm{\pi}}=-i\hbar\,\bm{\nabla}-q\,\mathbf{A}$ and $\hat{\mathbf{v}}=\hat{\bm{\pi}}/m$:
$\frac{d}{dt}\langle\hat{\mathbf{x}}\rangle=\frac{1}{m}\langle\hat{\bm{\pi}}\rangle,\qquad\frac{d}{dt}\langle\hat{\bm{\pi}}\rangle=-\langle\bm{\nabla}V\rangle+q\,\big\langle\mathbf{E}+\hat{\mathbf{v}}\times\mathbf{B}\big\rangle,$ (125)
@@ 7 Conclusion
Quantum kinematics and measurement descend from the same data. The states are rays in a Hermitian bundle with unitary connection $D$, and the unique probability assignment compatible with our axioms is the Born rule, fiberwise, its spacetime version follows by integrating the hh–norm on Cauchy slices with the induced measure $\sqrt{\gamma}\,d\Sigma$. This ties probabilities to the same real-slice geometry that governs dynamics.
On the perturbative side, implementing entire-function regulators at covariant kinetic operators and vertices yields a UV-finite framework, whose IR limit reproduces local EYM and matter. The unified Noether–II content includes covariant conservation laws and Slavnov–Taylor identities, ensuring consistency of the regulated theory.
@@
Several targeted directions can be follow, a group or metric locking principle either geometric or topological is needed to fix the relative normalization and uniquely determine the internal group, elevating our shared-structure unification toward a strict GUT selection. On the quantum–classical interface, the covariant Ehrenfest lemma, linking the time evolution of expectation values in quantum theory to the classical equations of motion, shows that geodesic drift and Lorentz force emerge in expectation directly from the unified geometry, extending this to open-system dynamics and decoherence on curved backgrounds.
Acknowledgments
We thank Martin Green and Arvin Kouroshnia for helpful discussions. Research at the Perimeter Institute for Theoretical Physics is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Research and Innovation (MRI).

### v2

@@ 6 Quantum Mechanics and Quantum Field Theory from HUFT
We use $\gamma^{a}$ ($a,b=0,1,2,3$) for the Dirac gamma matrices in a local Lorentz frame and:
$\gamma^{ab}:=\tfrac{1}{2}[\gamma^{a},\gamma^{b}]$ (130)
for their antisymmetrized products appearing in the spinor representation of $\mathfrak{spin}(1,3)$. These objects are not related to the induced spatial metric $\gamma_{ij}$ ($i,j=1,2,3$) used below in canonical or Schrödinger formulations. When we foliate spacetime by spacelike hypersurfaces $\Sigma_{t}$, the induced metric is $\gamma_{\mu\nu}=h_{\mu\nu}+n_{\mu}n_{\nu}$ with unit normal $n^{\mu}$, and $\gamma_{ij}$ denotes its pullback to coordinates on $\Sigma_{t}$; we write $\sqrt{\gamma}:=\sqrt{\det(\gamma_{ij})}$ for the corresponding volume density. We fix a spacelike hypersurface $\Sigma_{t}$ with induced metric $\gamma_{ij}$ and volume form $d\mu_{\Sigma_{t}}=\sqrt{\gamma}\,d^{3}x$. The one-particle Hilbert space is $L^{2}(\Sigma_{t},d\mu_{\Sigma_{t}})$ or its spinor analogue, with inner product:
$(\phi,\psi)_{\Sigma_{t}}:=\int_{\Sigma_{t}}\sqrt{\gamma}\,\phi^{*}\psi\,d^{3}x.$ (131)
A Schrödinger evolution is specified by an essentially self-adjoint Hamiltonian $\hat{H}$ on a dense domain, so that $i\partial_{t}\psi=\hat{H}\psi$ and probability conservation is $(\psi,\psi)_{\Sigma_{t}}=\mathrm{const}$.
We now will justify the appearance of the Schrödinger equation on the foliated real slice from the geometric probability structure already defined above. For probabilities on $\Sigma_{t}$ and norm conservation, given a state $\psi$ on $\Sigma_{t}$, the Born density is the fiberwise Hermitian norm integrated against the induced hypersurface measure $d\mu_{\Sigma_{t}}=\sqrt{\gamma}\,d^{3}x$, so the total probability on $\Sigma_{t}$ is:
$\|\psi\|^{2}_{\Sigma_{t}}:=(\psi,\psi)_{\Sigma_{t}}=\int_{\Sigma_{t}}\sqrt{\gamma}\,\psi^{\dagger}\psi\,d^{3}x.$ (132)
We impose the physical requirement that total probability is independent of the chosen Cauchy slice:
$\frac{d}{dt}\|\psi\|^{2}_{\Sigma_{t}}=0.$ (133)
This is the geometric content of probability conservation in the canonical picture. Next, identifying the Hilbert spaces at different $t$, because the inner product depends on the slice, we reduce to a fixed Hilbert space so we let $\Phi_{t}:\Sigma_{0}\to\Sigma_{t}$ be the diffeomorphism generated by the foliation flow such as the normal flow or equivalently, a choice of lapse/shift fixing an identification of points between slices. We write $J_{t}(x)$ for the Jacobian determinant relating the induced volume forms:
$(\Phi_{t})^{*}(\sqrt{\gamma(t)}\,d^{3}x)=J_{t}(x)\,\sqrt{\gamma(0)}\,d^{3}x.$ (134)
Define an isometry $I_{t}:\mathcal{H}_{t}\to\mathcal{H}_{0}$ by the weighted pullback:
$(I_{t}\psi_{t})(x):=J_{t}(x)^{1/2}\,\psi_{t}(\Phi_{t}(x)),$ (135)
so that for any $\phi_{t},\psi_{t}\in\mathcal{H}_{t}$:
$(\,I_{t}\phi_{t},\,I_{t}\psi_{t}\,)_{\Sigma_{0}}$ $=\int_{\Sigma_{0}}\sqrt{\gamma(0)}\,(I_{t}\phi_{t})^{\dagger}(I_{t}\psi_{t})\,d^{3}x$ (136)
$=\int_{\Sigma_{0}}\sqrt{\gamma(0)}\,J_{t}\,\phi_{t}(\Phi_{t}(x))^{\dagger}\psi_{t}(\Phi_{t}(x))\,d^{3}x$ (137)
$=\int_{\Sigma_{t}}\sqrt{\gamma(t)}\,\phi_{t}^{\dagger}\psi_{t}\,d^{3}x=(\phi_{t},\psi_{t})_{\Sigma_{t}}.$ (138)
Hence $I_{t}$ is unitary, an inner-product preserving identification. Unitary time evolution and the one-parameter group come when we let $U_{t\leftarrow 0}:\mathcal{H}_{0}\to\mathcal{H}_{t}$ denote the physical evolution map from $\Sigma_{0}$ to $\Sigma_{t}$, so that $\psi_{t}=U_{t\leftarrow 0}\psi_{0}$. Probability conservation (133) implies:
$\|\psi_{t}\|_{\Sigma_{t}}=\|\psi_{0}\|_{\Sigma_{0}}\quad\Longleftrightarrow\quad(U_{t\leftarrow 0}\phi_{0},\,U_{t\leftarrow 0}\psi_{0})_{\Sigma_{t}}=(\phi_{0},\psi_{0})_{\Sigma_{0}}.$ (139)
Equivalently, the map $\widetilde{U}(t):=I_{t}\circ U_{t\leftarrow 0}:\mathcal{H}_{0}\to\mathcal{H}_{0}$ is unitary:
$(\widetilde{U}(t)\phi_{0},\widetilde{U}(t)\psi_{0})_{\Sigma_{0}}=(\phi_{0},\psi_{0})_{\Sigma_{0}}.$ (140)
We assume the standard composition and continuity properties of time evolution:
$\widetilde{U}(0)=\mathbf{1},\qquad\widetilde{U}(t+s)=\widetilde{U}(t)\widetilde{U}(s),\qquad\widetilde{U}(t)\ \text{strongly continuous in }t.$ (141)
Then $\{\widetilde{U}(t)\}_{t\in\mathbb{R}}$ is a strongly continuous one-parameter unitary group on $\mathcal{H}_{0}$. The existence of a self-adjoint generator and the Schrödinger equation. By Stone’s theorem or more precisely the Stone–von Neumann theorem, there exists a densely-defined self-adjoint operator $\widehat{H}$ on $\mathcal{H}_{0}$ such that:
[…]
Differentiating (142) in the strong sense yields:
$\frac{d}{dt}\widetilde{U}(t)=-\frac{i}{\hbar}\,\widehat{H}\,\widetilde{U}(t).$ (143)
[…]
$i\hbar\,\partial_{t}\widetilde{\psi}(t)=\widehat{H}\,\widetilde{\psi}(t).$ (144)
Returning to the time-dependent slice using $\psi_{t}=I_{t}^{-1}\widetilde{\psi}(t)$ gives the canonical Schrödinger form on $\Sigma_{t}$:
$i\hbar\,\partial_{t}\psi_{t}=\widehat{H}_{t}\,\psi_{t},\qquad\widehat{H}_{t}:=I_{t}^{-1}\widehat{H}\,I_{t},$ (145)
with $\widehat{H}_{t}$ self-adjoint with respect to $(\cdot,\cdot)_{\Sigma_{t}}$. Thus, once the geometric probability assignment and norm conservation are imposed, the existence of a Schrödinger generator is not an additional assumption but follows from unitarity.
The real-slice geometry provides the induced spatial metric $\gamma_{ij}$ and the unitary covariant derivative $D_{\mu}$ including spin and internal gauge pieces. We denote by $D_{i}$ the pullback of $D_{\mu}$ to $\Sigma_{t}$ and define the spatial kinetic momentum operator:
$\widehat{\Pi}_{i}:=-i\hbar\,D_{i}.$ (146)
The natural gauge- and diffeomorphism-covariant Laplace–Beltrami operator on $\Sigma_{t}$ is:
$\Delta_{\gamma,D}\psi:=\frac{1}{\sqrt{\gamma}}\,D_{i}\!\left(\sqrt{\gamma}\,\gamma^{ij}D_{j}\psi\right),$ (147)
which is symmetric on $L^{2}(\Sigma_{t},d\mu_{\Sigma_{t}})$ under standard falloff/boundary conditions. For a spin-0 Schrödinger field with real scalar potential $V$, the minimally coupled Hamiltonian is:
$\widehat{H}_{t}=-\frac{\hbar^{2}}{2m}\,\Delta_{\gamma,D}+V,$ (148)
and for charged matter one includes the appropriate $U(1)$ or nonabelian temporal connection component inside $D_{t}$ (equivalently as a $qA_{0}$ term in $\widehat{H}_{t}$ depending on conventions). More generally, the Pauli/Dirac Hamiltonians are obtained by replacing (148) with the corresponding first-order operators built from $(h_{\mu\nu},D_{\mu})$, and the above unitary-generator argument applies verbatim.
Under these standard hypotheses self-adjointness and suitable boundary conditions, the Heisenberg identity:
$\frac{d}{dt}\langle\hat{A}\rangle=\frac{i}{\hbar}\langle[\hat{H},\hat{A}]\rangle+\left\langle\frac{\partial\hat{A}}{\partial t}\right\rangle$ (149)
@@
follows by differentiating $\langle\hat{A}\rangle$ and substituting the Schrödinger equation. We define the kinetic momentum operator $\hat{\Pi}_{\mu}:=-i\hbar D_{\mu}$, the position operator $\hat{X}^{\mu}$, multiplication by the coordinate function on $\Sigma_{t}$, and for any bundle endomorphism observable $\hat{O}$ define the expectation value on $\Sigma_{t}$ by:
@@
We take dynamics to be generated by the minimally coupled covariant Hamiltonian $\hat{H}$, either Dirac, Pauli, or Schrödinger built from $h$ and $D_{\mu}$, so that $i\hbar\,\partial_{t}\psi=\hat{H}\psi$ and the probability current is conserved.
We assume standard falloff, so boundary terms at spatial infinity vanish, and take $V$ a real scalar section such as a gauge singlet potential. Then, the expectation values of position and kinetic momentum obey:
@@
Conservation of the probability current gives the covariant continuity equation $\partial_{t}(\sqrt{\gamma}\,\psi^{\dagger}\psi)+\nabla_{i}(\sqrt{\gamma}\,j^{i})=0$, so boundary terms from $\partial_{t}\sqrt{\gamma}$ and $\nabla_{i}$ cancel in the time derivative of (150). Thus, we obtain the Heisenberg identity in covariant form:
@@
with the Levi–Civita correction precisely reproducing the $\Gamma$–terms below. For $\hat{O}=\hat{X}^{\mu}$, minimal coupling implies $\hat{\Pi}_{\alpha},\hat{X}^{\mu}]=-i\hbar\,\delta^{\mu}{}_{\alpha}$, and for the kinetic Hamiltonian $\hat{H}_{\rm kin}=\tfrac{1}{2m}\,h^{\alpha\beta}\hat{\Pi}_{\alpha}\hat{\Pi}_{\beta}$ we obtain
@@
yielding (151) after taking expectation values and accounting for the foliation, with no explicit time dependence of $\hat{X}^{\mu}$.
For $\hat{O}=\hat{\Pi}_{\mu}$, we use curvature and metric compatibility and define the kinetic momentum operator:
@@
meaning $h^{\alpha\beta}$ is covariantly constant, so it commutes with $\hat{\Pi}_{\mu}$ so there are no extra terms from derivatives of the metric when you commute things. Then the Heisenberg force operator follows:
@@
the right-hand side is the symmetrized product of momentum with field strength, so the operator is Hermitian. Its symmetric part gives the Lorentz-force term $q\,F_{\mu\nu}\,\dot{X}^{\nu}$ in expectation values. The scalar potential contributes $(i/\hbar)[V,\hat{\Pi}_{\mu}]=-\nabla_{\mu}V$. Finally, when transporting the expectation value with the time-dependent volume form and frame, the Levi–Civita piece from $D/dt$ supplies $\langle\Gamma^{\rho}_{\mu\nu}\,\hat{\Pi}_{\rho}\,\dot{X}^{\nu}\rangle$, completing (152). Since on the real slice, $B_{\mu\nu}=F_{\mu\nu}$, the gauge-force term is literally the antisymmetric part of the Hermitian metric.
In a local inertial frame at a point $\Gamma=0$, the relations are reduced by taking the expectation values and writing:
@@
When transporting the expectation value, that is the gravitational part. That extra term is what we call the Levi–Civita piece. It is the correction needed so that the equation of motion is tensorial coordinate-independent. Without it, you would be differentiating components as if the basis were fixed, flat space. With it, the quantum expectation values reduce to the geodesic with Lorentz force in the classical limit.
@@
On a static slice of flat space with coordinates $(t,\mathbf{x})$, $\Gamma=0$ and $h_{\mu\nu}\to\eta_{\mu\nu}$. Writing $\hat{\boldsymbol{\pi}}=-i\hbar\,\boldsymbol{\nabla}-q\,\mathbf{A}$ and $\hat{\mathbf{v}}=\hat{\boldsymbol{\pi}}/m$:
$\frac{d}{dt}\langle\hat{\mathbf{x}}\rangle=\frac{1}{m}\langle\hat{\boldsymbol{\pi}}\rangle,\qquad\frac{d}{dt}\langle\hat{\boldsymbol{\pi}}\rangle=-\langle\boldsymbol{\nabla}V\rangle+q\,\big\langle\mathbf{E}+\hat{\mathbf{v}}\times\mathbf{B}\big\rangle,$ (165)
@@ 7 Conclusions
Quantum kinematics and measurement descend from the same data. The states are rays in a Hermitian bundle with unitary connection $D$, and the unique probability assignment compatible with our axioms is the Born rule. Fiberwise, its spacetime version follows by integrating the $h$-norm on Cauchy slices with the induced measure $\sqrt{\gamma}\,d\Sigma$. This ties probabilities to the same real-slice geometry that governs dynamics.
On the perturbative side, implementing entire-function regulators at covariant kinetic operators and vertices yields a UV-finite framework, whose IR limit reproduces local EYM and matter. The unified Noether-II content includes covariant conservation laws and Slavnov–Taylor identities, ensuring consistency of the regulated theory.
@@
Several targeted directions can be followed. A group or metric locking principle, either geometric or topological, is needed to fix the relative normalization and uniquely determine the internal group, elevating our shared-structure unification toward a strict GUT selection. On the quantum–classical interface, the covariant Ehrenfest lemma, linking the time evolution of expectation values in quantum theory to the classical equations of motion, shows that geodesic drift and Lorentz force emerge in expectation directly from the unified geometry, extending this to open-system dynamics and decoherence on curved backgrounds.
[…]
Conceptualization, E.J.T. and J.W.M.; methodology, E.J.T.; investigation, E.J.T. and J.W.M.; writing—original draft preparation, E.J.T.; writing—review and editing, E.J.T. and J.W.M.; supervision, J.W.M. All authors have read and agreed to the published version of the manuscript.
[…]
This research received no external funding.
[…]
No new data were created or analyzed in this study. Data sharing is not applicable.
Acknowledgements.
[…]